# Response Tuning: Aligning Large Language Models without Instruction

### v1

@@ Abstract
Instruction tuning—supervised fine-tuning using instruction-response pairs—is a foundational step in transitioning pre-trained Large Language Models (LLMs) into helpful and safe chat assistants. Our hypothesis is that establishing an adequate output space can enable such a transition given the capabilities inherent in pre-trained LLMs. To verify this, we propose Response Tuning (RT), which eliminates the instruction-conditioning step in instruction tuning and solely focuses on response space supervision. Our experiments demonstrate that RT models, trained only using responses, can effectively respond to a wide range of instructions and exhibit helpfulness comparable to that of their instruction-tuned counterparts. Furthermore, we observe that controlling the training response distribution can significantly improve their user preference or elicit target behaviors such as refusing assistance for unsafe queries. Our findings illuminate the role of establishing an adequate output space in alignment, highlighting the potential of the extensive inherent capabilities of pre-trained LLMs. [^1] Codes are available at https://github.com/seokhyunan/response-tuning.
@@ 1 Introduction
Large Language Models (LLMs) are pre-trained to predict the next token using massive amounts of web-crawled text, implicitly learning a wide range of tasks (Radford et al., 2019; OpenAI, 2023; Dubey et al., 2024). To align them with human needs, LLMs typically undergo Instruction Tuning (IT) (Mishra et al., 2022; Wei et al., 2022a; Sanh et al., 2022)—a supervised fine-tuning process using instruction-response paired data. This conditional supervision enables them to follow natural language instructions and responsibly handle unsafe queries, facilitating their real-world applications (Wang et al., 2023a; b; Ivison et al., 2023; OpenAI, 2023; Xu et al., 2024; Zhou et al., 2024; Bianchi et al., 2024; Dubey et al., 2024).
We note that LLMs might acquire abilities such as instruction-following and safety assessment during pre-training, as well as the extensive knowledge required to perform specific tasks (Radford et al., 2019; Brown et al., 2020; Zhou et al., 2024). Moreover, previous observations suggest that output space information plays a crucial role in canonical NLP tasks (Min et al., 2022; Kung & Peng, 2023). Our hypothesis is that establishing an adequate output space can surface these capabilities in the form of responses, enabling pre-trained LLMs to function as desired chat assistants.
In this study, we explore the impact of establishing response space in transforming pre-trained LLMs into helpful and safe assistants. To this end, we propose Response Tuning (RT), which focuses solely on response space supervision. Specifically, RT omits the instruction-conditioning process and the model is not supervised from instruction-response mappings (see Figure 1). Instead, it trains the model to construct responses and learn their distribution. This omission allows us to examine the isolated effect of response supervision.
We first investigate whether establishing an output space alone can enable LLMs to generate aligned responses without additional supervision from instruction-response mappings. To this end, we evaluate the instructability of RT models trained on four recent LLMs using three different datasets: Llama-3.1-8B (Dubey et al., 2024), Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023), utilizing only the responses from Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), and LIMA (Zhou et al., 2024). Our human and automatic evaluations, based on test instructions derived from multiple sources (Li et al., 2023; Lin et al., 2024), show that RT models can appropriately respond to a wide range of instructions and exhibit preference comparable to their IT counterparts. They also demonstrate core capabilities (reasoning, knowledge, etc.) on par with those of their counterparts. These findings suggest that instruction-following capabilities are largely acquired during pre-training and that establishing an appropriate response space can yield instruction-following assistants.
@@
Figure 1: Comparison of IT and RT. In both methods, the loss is computed exclusively on the response tokens. Unlike IT, RT does not involve instruction conditioning and focuses solely on learning the distribution of responses.
Furthermore, we demonstrate that controlling the training response distribution can achieve the target alignment. Specifically, we observe that refining the structural attributes of the training responses leads to significant improvements in user preference, enabling the models to assist users more effectively. Moreover, we find that incorporating a small set of contextual refusals—which embed guidance for handling unsafe queries—into the RT data allows RT models to implicitly evaluate and reject unsafe queries. When tested using multiple safety benchmarks (Zou et al., 2023; Huang et al., 2024; Mazeika et al., 2024; Röttger et al., 2024), these models exhibit refusal rates comparable to IT counterparts trained with paired examples. Additionally, these phenomena become more pronounced as the base model’s capabilities scale. These findings suggest that strong base LLMs can acquire complex capabilities during pre-training and that adequately embedding behavioral guidance into the response space can invoke target behaviors.
@@
Finally, we show that our observations still hold in an in-context learning setting. We demonstrate that in-context learning with response demonstrations only can yield helpful and safe assistants, further supporting our hypothesis. Our work highlights the impact of establishing a desired response distribution in LLM alignment and emphasizes the extensive inherent capabilities of pre-trained LLMs.
@@
    We hypothesize that establishing an appropriate output space may enable pre-trained LLMs to become instruction-following and safe assistants. To verify this, we propose Response Tuning (RT), which does not involve instruction-conditioning and focuses on learning the response distribution.
@@
    Our extensive evaluations show that RT models, trained solely using responses, can effectively respond to a wide range of instructions. This indicates that instruction-following capabilities are largely acquired during pre-training and establishing an output space can yield instruction-following assistants.
@@
    Furthermore, we demonstrate that the target alignment can be achieved by controlling the training response distribution. Notably, we demonstrate that refining the structural attributes of responses and incorporating a small number of contextual refusals can improve user preference and safety, respectively.
@@ 2 Related Work
Instruction tuning. Instruction Tuning (IT) is a process of supervised fine-tuning LLMs using instruction-response pairs, where the model is trained to produce responses (assistant outputs) conditioned on paired instructions (user inputs). Recent studies, expanding from earlier research that focused on cross-task generalization across canonical NLP tasks (Mishra et al., 2022; Wei et al., 2022a; Sanh et al., 2022), have shifted their focus to generalization for unseen user instructions, demonstrating remarkable success in improving the usability of LLMs. Notable contributions include synthetic data generation frameworks (Wang et al., 2023b; Honovich et al., 2023; Ding et al., 2023; Xu et al., 2024), human-involved conversation collection schemes (Conover et al., 2023; Köpf et al., 2023; Chiang et al., 2023; Zhou et al., 2024; Zhao et al., 2024; Zheng et al., 2024), and efficient training techniques (Zhou et al., 2024; Lin et al., 2024; Chen et al., 2024; Liu et al., 2024). However, few controlled studies have investigated the impact of specific learning signals on producing instruction-following assistants.
Large Language Model safety. Instruction-following LLMs are susceptible to malicious use without proper safeguards, and their risks become more nuanced and pronounced as their capabilities grow (Bommasani et al., 2021; Hendrycks et al., 2023; Kang et al., 2023). This has motivated studies on safety measures to mitigate their potential risks in real-world applications (Perez et al., 2022; Bai et al., 2022b; Dubey et al., 2024; Mu et al., 2024). These measures typically aim to make the model refuse to answer unsafe queries—such as requests for assistance in committing illegal activities—while retaining their helpfulness in appropriate contexts. A common practice to achieve this is incorporating unsafe instruction-refusal response pairs into the IT stage (Touvron et al., 2023; Bianchi et al., 2024; Dubey et al., 2024). However, as with instruction-following alignment, it remains unclear which cues mainly elicit the desired behaviors.
Supervision from input-output pairs. Min et al. (2022) show that the input-output mapping matters significantly less than the label space information in the in-context demonstrations of canonical NLP tasks (e.g., classification). Taking a step further, later works examine the role of task definitions given in the prompts of the IT dataset and demonstrate that models trained with minimal (only label space) or misleading task definitions have little effect on their task generalization performance (Kung & Peng, 2023). However, those findings are limited to conventional label spaces, which correspond to predefined categories or outputs expected in given NLP tasks (e.g., yes/no, positive/negative), rather than the open-ended output space in open-domain IT.
Our approach. To address these gaps, we aim to examine the impact of supervision from the response distribution in alignment. Specifically, we focus on open-domain, open-ended IT—a widely adopted practice in aligning both open-source and proprietary LLMs (Wang et al., 2023a; Köpf et al., 2023; Jiang et al., 2023; OpenAI, 2023; Dubey et al., 2024; Riviere et al., 2024). Our focus lies on two foundational objectives of LLM alignment: transforming pre-trained LLMs into instruction-following and safe assistants.
@@ 3 Response Tuning (RT)
To verify our hypothesis, we propose Response Tuning (RT) to examine the impact of establishing a response space in alignment.
Training data. We adopt the chatbot-style schema proposed by Wang et al. (2023a), which separates user instructions and assistant responses using special tokens: <|user|> and <|assistant|>. However, in RT, we omit both the <|user|> token and the user instructions during training. Therefore, the training data consists only of the <|assistant|> token followed by the training response.
@@
Here, $\theta$ represents the model parameters, $l$ is the total number of tokens in a training example, and $t_{i}$ denotes the $i$-th token in the sequence. The indicator function $\delta_{i}$ equals 1 if the $i$-th token belongs to the response set $\mathbf{R}$ (i.e., the assistant’s response), and 0 otherwise. This formulation mirrors the loss function used in IT, where loss masking is applied to instruction tokens (Wang et al., 2023b; Sanh et al., 2022; Wang et al., 2023a; Dettmers et al., 2024). However, unlike IT, RT does not condition the response tokens on the paired instructions, allowing the model to focus on learning the distribution of responses independently.
Inference. We input a sequence that starts with the instruction delimiter (<|user|>), followed by the user’s instruction, and then the response delimiter (<|assistant|>). The model then generates the assistant’s response after the <|assistant|> token. Although the RT models are trained without paired instructions, they can leverage their inherent capabilities to generate desired responses. We will demonstrate this in the subsequent experiment sections.
@@ 4 Instructability of RT Models
In this section, we evaluate the instructability of RT models. We first assess their open-ended instruction-following capabilities to determine whether they can handle typical user queries in real-world applications (e.g., creative writing, trip planning, and general question-answering). Additionally, we test their core capabilities needed to perform instructed tasks using multiple benchmarks. For reliable verification, we conduct both human and automatic evaluations.
@@ 4.1 Experimental Setup
Pre-trained LLMs. We use four recent open-source LLMs: Llama-3.1-8B (Dubey et al., 2024), Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023). These models are widely adopted due to their high performance, steerability, and efficiency. In this section, we mainly report results for Llama-3.1-8B and Gemma-2-9B; results for the other models are provided in Appendix C.
Training dataset. To account for potential dataset dependency of RT, we use three different IT datasets, from which we only use the response subsets:
  • Alpaca (Taori et al., 2023): A dataset of 52,000 instruction-response pairs generated through the Self-Instruct(Wang et al., 2023b) framework. We use its cleaned version, which fixes issues in the original dataset with the updated GPT-4 (OpenAI, 2023) generation backbone. [^2] https://huggingface.co/datasets/yahma/alpaca-cleaned
  • Dolly (Conover et al., 2023): A dataset of 15,000 instruction-response pairs manually crafted by human annotators.
  • LIMA (Zhou et al., 2024): A dataset of 1,000 instruction-response pairs curated from various sources, including web data from Stack Exchange, wikiHow, and Reddit, as well as examples manually written by the authors and sourced from Super-NaturalInstructions (Wang et al., 2022).
Training setup. We use a parameter-efficient fine-tuning method, QLoRA (Dettmers et al., 2024), which has been shown to match the performance of full 16-bit fine-tuning while significantly reducing memory footprint. LoRA adapters (Hu et al., 2022) are applied to all linear layers and are double-quantized in 4-bit NormalFloat during training. We set the alpha, rank, and dropout rate of the adapters to 16, 64, and 0.1, respectively. A 32-bit paged AdamW optimizer (Dettmers et al., 2024) is used with a batch size of 64 and a constant learning rate of 1e-4 (Wang et al., 2022; Wei et al., 2022a). Models are trained for 10 epochs with a maximum token length of 2,048 using NVIDIA A6000 (48GB VRAM) or A100 (80GB VRAM) GPUs. We use vLLM with greedy decoding for generation (Kwon et al., 2023). This setup is applied to all experiments unless otherwise specified.
Instructabiltiy evaluation. To determine whether RT models can handle user instructions, we independently assess their responses’ acceptability and quality. Additionally, we conduct pairwise assessments to measure their preference compared to their IT counterparts.
@@
Figure 2: Human evaluation of response acceptability for RT and IT models. Evaluators rate responses to 805 test instructions as ‘Acceptable (Excellent)’, ‘Acceptable (Sufficient)’, or ‘Not Acceptable’. The results indicate that RT models can appropriately respond to a broad range of instructions.
@@
Figure 3: GPT-4 evaluation of response quality for RT and IT models. The test is conducted using 800 regular instructions from the JustEval benchmark (Lin et al., 2024). The results indicate that RT models achieve performance close to IT models across all metrics.
@@
Figure 4: Pairwise preference evaluation results for Llama-3.1-8B-based RT models. Human evaluators and the GPT-4 judge are asked to choose the more helpful response between RT and IT models for the same instruction. The results indicate that RT models exhibit preferences comparable to their IT counterparts.
Table 1: Evaluation of core capabilities for IT and RT models. RT models exhibit similar results to IT models across all benchmarks used. This finding indicates that RT models retain the core capabilities required to assist users.
@@
RT 56.87 32.20 56.23 60.55 43.59 74.86 58.42
Gemma-2-9B $+$ Alpaca IT 65.22 39.00 52.68 61.33 67.78 76.88 64.78
RT 64.35 38.40 59.29 61.67 66.41 76.39 65.62
@@
Response acceptance rate ($\uparrow$) 0.84 0.90
LC Win-rate against IT models ($\uparrow$) 37.58 40.00
Table 3: Preference evaluation results for instruction-following Llama-3.1-8B models trained using refined responses. The models are compared with those trained on the original datasets. Except for Alpaca (GPT-4 generated), models trained on refined responses are significantly preferred over their counterparts.
Method Dataset LC Win Rate (%)
(vs non-refined)
Alpaca (response refined) 53.37
IT Dolly (response refined) 68.75
LIMA (response refined) 62.88
Alpaca (response refined) 49.39
RT Dolly (response refined) 66.09
LIMA (response refined) 63.48
  • Independent Assessment: To evaluate the acceptability of open-ended responses from RT models, we conduct a human evaluation. We use the test instructions from AlpacaEval (Li et al., 2023), a mixture of five distinct evaluation datasets covering a wide range of instructions (Wang et al., 2023b; Köpf et al., 2023; Bai et al., 2022a; Chiang et al., 2023; Geng et al., 2023). Evaluators are presented with an instruction and the corresponding model response, and are asked to rate the responses by choosing one of three options: Acceptable (Excellent), Acceptable (Sufficient), or Not Acceptable. To address the limitations of human evaluators (Gudibande et al., 2024), we also conduct fine-grained response quality evaluation using the JustEval benchmark (Lin et al., 2024), which utilizes a GPT-4 judge. The evaluation interface, guidelines, and prompts can be found in Appendix A.
  • Pairwise Assessment: We conduct both human and automatic evaluations using the AlpacaEval test instructions. Human evaluators are asked to choose the more helpful response or indicate a Tie, given the instruction and responses from both the RT model and its IT counterpart. For automatic evaluation, we utilize the GPT-4 judge of AlpacaEval 2.0 (Li et al., 2023) and report length-controlled win rates (Dubois et al., 2024). This evaluator exhibits a high Spearman correlation (Spearman, 1904) of 0.98 with human evaluators in Chatbot Arena (Chiang et al., 2024).
Core capabilities evaluation. Recent work argues that a general-purpose assistant should be able to perform core tasks to ensure practicality (Wang et al., 2023a). To measure the core capabilities of RT models, we conduct automated evaluations using multiple benchmarks designed to assess specific skills. The following benchmarks are considered: MMLU (Hendrycks et al., 2021) and OpenbookQA (Mihaylov et al., 2018) for knowledge, HellaSwag (Zellers et al., 2019) for commonsense, ARC (Clark et al., 2018) for reasoning, GSM8K (Cobbe et al., 2021) for mathematical reasoning, and PIQA (Bisk et al., 2020) for physical reasoning.
4.2 Results
Establishing a response space alone enables LLMs to behave as chat assistants. Figures 23 and 4 depict the human acceptability, response quality, and pairwise helpfulness assessment results, respectively. The independent assessment results indicate that RT models, trained only using responses without paired instructions, are capable of generating appropriate responses to a wide range of instructions. A significant majority of the responses generated by RT models are rated as Acceptable, with many achieving the Excellent rating. They perform comparably to their IT counterparts in response quality evaluations, achieving similar scores across all metrics of the JustEval benchmark. The pairwise evaluation results also show that RT models exhibit competent, though slightly inferior, preference compared to their IT counterparts. This trend remains consistent across four different base LLMs and three different datasets. These findings suggest that the instruction-response dynamics are largely acquired during pre-training, and just establishing a response space can elicit those dynamics to generate aligned responses. The examples of the model outputs are available in Appendix E.
RT models exhibit core capabilities on par with IT models. Table 1 displays the core capabilities evaluation results. While there exist some fluctuations due to the nature of fine-tuning (Fu et al., 2024), we find no significant differences between the performance of RT and IT models across all tasks tested. These results indicate that RT models can comprehend inputs, adhere to constraints to yield answers in a zero-shot setting, and possess core capabilities comparable to IT models. These core capabilities that RT models retain from pre-trained models can serve as the foundation in assisting users with diverse requests. Additionally, it supports previous findings that input-output mappings do not significantly impact end performance and that most knowledge required for specific tasks is acquired during pre-training (Kung & Peng, 2023; Zhou et al., 2024).
Instruction-following capabilities are better internalized in stronger LLMs. Table 3 shows the evaluation results for RT Gemma-2-2B and Gemma-2-9B models. While smaller models like Gemma-2-2B generate highly acceptable responses, larger models such as Gemma-2-9B, which are pre-trained on significantly more tokens (2T vs. 8T tokens), exhibit higher human acceptance rates and preferences against IT counterparts. This suggests that stronger models are better at learning instruction-following capabilities during pre-training.
5 Refining Response Distribution for Preference Alignment
5.1 Refining Response Attributes
Motivation. Our preceding experiments demonstrate that establishing a response space alone can enable LLMs to behave as instruction-following assistants. In this section, we investigate whether refining the training response distribution can further enhance user preference, which is one of the essential objectives of LLM alignment. While previous works have shown that techniques like feedback learning or completely regenerating responses in the IT dataset using other aligned LLMs can improve user preference (Bai et al., 2022a; Peng et al., 2023; Ivison et al., 2024), we are interested in verifying whether adjusting the inherent attributes of existing response distributions in IT or RT data can yield similar benefits.
5.2 Experimental Setup
Response refinement. We focus on refining the inherent attributes of the response space, specifically considering three core elements that correlate with user preference: clarity, structure, and tone. For the refinement, we utilize a strong instruction-following LLM, Llama-3.1-70B-Instruct (Dubey et al., 2024), with a manually designed refinement prompt. The responses from the Alpaca, Dolly, and LIMA datasets are refined according to the following guidelines:
  • Clarity: Make the response easy to understand. It should be direct and to the point, avoiding complex language that might confuse the user
The prompt and the example of refined data are available in Appendix A and D.
Evaluation. We conduct simulated preference evaluations using the GPT-4 judge of AlpacaEval (Li et al., 2023) and report length-controlled win rates (Dubois et al., 2024). To directly measure the gains from response space refinement, we compare both the IT and RT models trained using the refined datasets with those trained on the original datasets.
5.3 Results
Refined response distribution yields models with better user preference. As shown in Table 3, the pairwise evaluation results indicate that both IT and RT models trained with the refined responses significantly outperform their counterparts trained on the original responses. The improvements are consistent across different base models and datasets, except for Alpaca. This exception may be due to Alpaca’s responses having limited room for improvement, as they are generated using GPT-4, a highly-aligned model using human feedback. These findings suggest that the training response distribution substantially contributes to the model’s ability to generate preferable and effective outputs, emphasizing the importance of training response distribution in achieving the alignment. The examples of model output can be found in Appendix E.

### v2

# Revealing the Inherent Instructability of Pre-Trained Language Models

@@ Abstract
Instruction tuning—supervised fine-tuning using instruction-response pairs—is a key step in making pre-trained large language models (LLMs) instructable. Meanwhile, LLMs perform multitask learning during their pre-training, acquiring extensive knowledge and capabilities. We hypothesize that the pre-training stage can enable them to develop the ability to comprehend and address instructions. To verify this, we propose Response Tuning (RT), which removes the instruction and its corresponding mapping to the response from instruction tuning. Instead, it focuses solely on establishing the response distribution. Our experiments demonstrate that RT models, trained only on responses, can effectively respond to a wide range of instructions and exhibit helpfulness approaching that of their instruction-tuned counterparts. In addition, we observe that the models can recognize and reject unsafe queries after learning the refusal conditions from training responses. Furthermore, we demonstrate that these observations also hold in an in-context learning setting. These findings support our hypothesis, highlighting the extensive inherent capabilities of pre-trained LLMs. [^1] Our codes are available at https://github.com/seokhyunan/response-tuning.
Revealing the Inherent Instructability of Pre-Trained Language Models
Seokhyun An1   Minji Kim2thanks: Work was partially conducted at UNIST.   Hyounghun Kim2footnotemark: 1Ulsan National Institute of Science and Technology (UNIST) 2Pohang University of Science and Technology (POSTECH) seokhyun@unist.ac.kr   {mzkim, imhkim}@postech.ac.kr
@@ 1 Introduction
Large Language Models (LLMs) are pre-trained to predict the next token using massive amounts of web-crawled text, implicitly learning a wide range of tasks (Radford et al., 2019; OpenAI, 2023; Dubey et al., 2024). To make the pre-trained models instructable, Instruction Tuning (IT) (Mishra et al., 2022; Wei et al., 2022a; Sanh et al., 2022)—a supervised fine-tuning process using instruction-response paired data—is widely performed, thereby enhancing their usability and facilitating real-world applications (Wang et al., 2023a, b; Ivison et al., 2023; OpenAI, 2023; Xu et al., 2024; Zhou et al., 2024; Bianchi et al., 2024; Dubey et al., 2024). However, how LLMs achieve such a transition remains unclear (Kung and Peng, 2023; Ghosh et al., 2024).
@@
Figure 1: Comparison of IT and RT. In both methods, the loss is computed exclusively on the response tokens. Unlike IT, RT omits the instruction and its corresponding mapping to the response, focusing solely on learning the response distribution.
We hypothesize that LLMs can acquire the ability to process instructions—that is, to comprehend and address them—during pre-training, in addition to the extensive knowledge required to perform specific tasks (Radford et al., 2019; Brown et al., 2020; Zhou et al., 2024). For instance, they may implicitly learn instruction-response dynamics from Q&A threads on Stack Exchange—a common source in typical pre-training datasets (Elazar et al., 2024). Under this assumption, the models may be able to respond appropriately to instructions once their response distribution is established. Previous studies have reported similar phenomena in canonical NLP tasks, demonstrating that supervision from the output space is crucial for performing such tasks (Min et al., 2022; Kung and Peng, 2023).
To test our hypothesis, we propose Response Tuning (RT), which omits the instruction and its associated instruction-response mapping from IT (see Figure 1). This omission prevents the model from learning to produce responses according to instructions; rather, it focuses on establishing the response distribution.
We evaluate whether RT models can understand and respond appropriately to diverse instructions. In our experiments, we consider four recent LLMs and three different datasets: Llama-3.1-8B (Dubey et al., 2024), Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023), utilizing only the responses from Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), and LIMA (Zhou et al., 2024). Our human and automatic evaluations, based on test instructions derived from multiple sources (Li et al., 2023; Lin et al., 2024), show that RT models can appropriately respond to a wide range of instructions and exhibit helpfulness close to that of their IT counterparts. These findings show that establishing a response distribution alone can yield instruction-following models, suggesting that the ability to process instructions is acquired during pre-training.
Furthermore, we find that RT models can identify and reject unsafe requests by invoking learned refusal conditions. Specifically, we incorporate explanatory refusals for unsafe queries—responses that decline to fulfill requests while specifying the underlying safety policy—into the training data. When tested with diverse unsafe instructions derived from multiple safety evaluation datasets (Zou et al., 2023; Huang et al., 2024; Mazeika et al., 2024; Röttger et al., 2024), these models exhibit refusal rates approaching those of IT counterparts trained with paired examples. These results also support our hypothesis that pre-training enables the model to develop the ability to process instructions.
@@
Finally, we show that our observations hold in an in-context learning setting. We demonstrate that in-context learning with response-only demonstrations can enable the models to appropriately handle instructions, further supporting our hypothesis.
@@
    We hypothesize that LLMs might acquire the ability to process instructions during pre-training. To verify this, we propose Response Tuning (RT), which omits both the instruction and its corresponding mapping to the response, focusing solely on establishing a response distribution.
@@
    Our extensive evaluations show that RT models—trained solely on responses—can effectively respond to a wide range of instructions, indicating that the ability to process instructions is acquired during pre-training.
@@
    Furthermore, we observe that they can recognize and reject unsafe requests by comprehending instructions and invoking learned refusal conditions. These observations are consistently replicated in an in-context learning setting, further supporting our hypothesis.
@@ 2 Related Work
Instruction tuning. Instruction Tuning (IT) is a supervised fine-tuning process using instruction-response paired data, where the model is trained to produce responses according to the instructions. Recent studies have shifted their focus from cross-task generalization in canonical NLP tasks (Weller et al., 2020; Mishra et al., 2022; Wei et al., 2022a; Sanh et al., 2022) to generalization for unseen user instructions. Notable contributions include synthetic data generation frameworks (Wang et al., 2023b; Honovich et al., 2023; Ding et al., 2023; Xu et al., 2024), human-involved conversation collection methods (Conover et al., 2023; Köpf et al., 2023; Chiang et al., 2023; Zhou et al., 2024; Zhao et al., 2024; Zheng et al., 2024). Recently, a line of research has explored the superficiality of the IT stage, questioning the necessity of large-scale data (Zhou et al., 2024; Chen et al., 2024; Liu et al., 2024) or parameter updates (Lin et al., 2024). However, what the models learn from IT remains unclear (Kung and Peng, 2023; Ghosh et al., 2024), and ablation studies exploring these aspects are limited.
Learning from input-output pairs. Min et al. (2022) highlights the crucial role of label space information in the in-context learning of canonical NLP tasks (e.g., classification). Kung and Peng (2023) examines the role of task definitions in the prompts of the IT dataset and demonstrates that models trained with misleading task definitions or only on the label space exhibit similar task generalization performance. However, these studies focus on generalization across canonical tasks with a finite label space, rather than on how models handle natural language instructions with an open-ended response space.
Our approach. To address these gaps, we investigate whether RT models—which have not been explicitly trained to process instructions—can understand and respond to them appropriately. A concurrent study by Hewitt et al. (2024) also examines whether training only on responses can yield instruction-following. In our work, we carefully investigate the inherent ability to process instructions using a multi-pronged evaluation, which includes refusal tests for unsafe queries and assessments of in-context response learning.
@@ 3 Response Tuning (RT)
We propose Response Tuning (RT) to verify our hypothesis that the ability to process instructions can be developed in the pre-training stage.
Training data. We adopt the chatbot-style template proposed by Wang et al. (2023a), which separates user instructions and assistant responses using special tokens: <|user|> and <|assistant|>. However, in RT, we omit both the <|user|> token and the instruction during training. Therefore, the training data consists only of the <|assistant|> token followed by the training response.
@@
Here, $\theta$ represents the model parameters, $l$ is the total number of tokens in a training example, and $t_{i}$ denotes the $i$-th token in the sequence. The indicator function $\delta_{i}$ equals 1 if the $i$-th token belongs to the response set $\mathbf{R}$ (i.e., the assistant’s response), and 0 otherwise. This formulation mirrors the loss function used in IT, where loss masking is applied to instruction tokens (Wang et al., 2023b; Sanh et al., 2022; Wang et al., 2023a; Dettmers et al., 2024). However, unlike IT, RT does not condition the response tokens on the paired instructions, which precludes the model from learning to generate responses according to instructions. Rather, it focuses on learning the response distribution.
Inference. We input a sequence that starts with the instruction delimiter (<|user|>), followed by the user’s instruction, and then the response delimiter (<|assistant|>). The model then generates the assistant’s response after the <|assistant|> token. Although RT does not involve explicitly learning how to process instructions, we find that the models can effectively leverage their inherent capabilities to do so. We demonstrate this in the following experimental sections.
@@ 4 Instructability of RT Models
In this section, we assess whether RT models can process diverse instructions (e.g., creative writing, trip planning, and general question-answering) and, if so, how their performance compares to that of IT models. For reliable verification, we conduct both human and automatic evaluations across four models and three datasets.
@@ 4.1 Experimental Setup
Pre-trained LLMs. We use four widely adopted recent LLMs: Llama-3.1-8B (Dubey et al., 2024), Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023). In this section, we mainly report results for Llama-3.1-8B and Gemma-2-9B; results for the other models are provided in Appendix C.
Training dataset. We use three different IT datasets, from which we only use the response subsets: [^2] Theoretically, RT can also be performed on general texts that lack paired instructions. We explore this approach in Appendix F.
@@
Figure 2: Human evaluation of response acceptability for RT and IT models. Evaluators rate responses to 805 test instructions as ‘Acceptable (Excellent)’, ‘Acceptable (Sufficient)’, or ‘Not Acceptable’. The results indicate that RT models can appropriately respond to diverse instructions. Refer to Table 8 for the results of the other models.
@@
Figure 3: GPT-4 response quality evaluation results for RT and IT models. We evaluate responses to 800 regular instructions from the JustEval benchmark (Lin et al., 2024) using GPT-4. The radar plots show the average GPT-4 ratings for each criterion. The results suggest that RT models achieve performance close to that of IT models on all metrics. Additional results for other models are available in Table 10.
@@
Figure 4: Pairwise evaluation results for Llama-3.1-8B-based RT models. Human evaluators and the GPT-4 judge are asked to choose the more helpful response between RT and IT models for the same instruction. The results indicate that RT models exhibit preferences similar to their IT counterparts. The results for the other models are provided in Tables 9 and 10.
@@
RT 56.87 32.20 56.23 60.55 43.59 74.86 54.05
- Untuned 63.36 33.6 60.04 66.34 55.72 80.14 59.87
Gemma-2-9B $+$ Alpaca IT 65.22 39.00 52.68 61.33 67.78 76.88 60.48
RT 64.35 38.40 59.29 61.67 66.41 76.39 61.08
Untuned 69.04 33.8 61.09 74.42 69.9 81.28 64.92
Table 1: Core capabilities evaluation results for IT and RT models. The results indicate that RT models largely retain their core capabilities and exhibit performance similar to that of IT models across all benchmarks. See Table 9 for results on the other models.
@@
Response acceptance rate 0.84 0.90
LC win-rate against IT models 37.58 40.00